OpenStudy (anonymous):

Find the quotient and remainder using synthetic division for \frac(x^3 + 3 x^2 + 4 x + 8)(x+2)

jimthompson5910 (jim_thompson5910):

Any ideas on how to get started?

OpenStudy (anonymous):

I'm pretty sure it's divison. that's why it says \frac in front.

jimthompson5910 (jim_thompson5910):

I know, the \frac means "fraction" but do you have any ideas on how to start the synthetic division?

OpenStudy (anonymous):

I have no idea what synthetic division even means.

jimthompson5910 (jim_thompson5910):

Have you learned polynomial long division?

OpenStudy (anonymous):

Yes, I'm doing it right now on a different problem but I don't know what synthetic divison is.

jimthompson5910 (jim_thompson5910):

synthetic division is like polynomial division, but you only focus on the coefficients and it's slightly faster and easier to remember (in my opinion)

jimthompson5910 (jim_thompson5910):

however, synthetic division can't be used all the time

OpenStudy (anonymous):

does it end up having a remainder at the end? If so, then I'm doing this current problem wrong and I've been doing this problem exactly right.

jimthompson5910 (jim_thompson5910):

to get started, you need to identify the coefficients x^3 + 3 x^2 + 4 x + 8 is the same as 1x^3 + 3 x^2 + 4 x + 8 The coefficients are 1, 3, 4, 8 Because we're dividing by x+2, this means we're using -2 as a test zero (since -2 makes x+2 equal to 0) So set up the synthetic division table to get -2 | 1 3 4 8 | ---------------------------------------- Then drop the first coefficient (1) to get this -2 | 1 3 4 8 | ---------------------------------------- 1 Then multiply the outer test zero (-2) by that number in the very bottom to get -2*1 = -2. Write this under the 3 like so -2 | 1 3 4 8 | -2 ---------------------------------------- 1 Then add 3 and -2 to get 3+(-2) = 1. This goes right under the -2 to get this -2 | 1 3 4 8 | -2 ---------------------------------------- 1 1 Then you repeat the last two steps -2 | 1 3 4 8 | -2 -2 ---------------------------------------- 1 1 -2 | 1 3 4 8 | -2 -2 ---------------------------------------- 1 1 2 Keep going until you fill up the second and third rows with numbers.

jimthompson5910 (jim_thompson5910):

So you need to do 3 more steps (because you'll stop when the last row has 4 values in it.

OpenStudy (anonymous):

I am so confused by this

jimthompson5910 (jim_thompson5910):

Do you have any notes about synthetic division? If not, then you may need to review it.

OpenStudy (anonymous):

No notes, nothing. My teacher just kind of shows things on the board and says "do this"

jimthompson5910 (jim_thompson5910):

Do my steps help at all?

OpenStudy (anonymous):

not really /:

jimthompson5910 (jim_thompson5910):

do you see how I got the coefficients 1 3 4 8

OpenStudy (anonymous):

yes I understand that. the numbers that are part of the problem are the coefficients, but I don't understand is how to exactly get the answer that I'm looking for

jimthompson5910 (jim_thompson5910):

do you see how I got the -2 (that's on the very left)

OpenStudy (anonymous):

basically plugging it in for x to make 0 or something like that

jimthompson5910 (jim_thompson5910):

yes, solving x+2 = 0 gives you x = -2

jimthompson5910 (jim_thompson5910):

so x+2 = 0 -2+2 = 0 0 = 0

jimthompson5910 (jim_thompson5910):

This gives us the starting table -2 | 1 3 4 8 | ----------------------------------------

jimthompson5910 (jim_thompson5910):

The next step after this is to write the first coefficient 1 under the line to get this -2 | 1 3 4 8 | ---------------------------------------- 1

jimthompson5910 (jim_thompson5910):

Step 3) Multiply -2 by 1 (the 1 that's in the bottom row) to get -2. Then write this under the 3 in the first row to get this -2 | 1 3 4 8 | -2 ---------------------------------------- 1 Step 4) Add 3 and -2 to get 1. Write this under the -2 -2 | 1 3 4 8 | -2 ---------------------------------------- 1 1

jimthompson5910 (jim_thompson5910):

You keep repeating these two steps until you reach the end of the row

OpenStudy (anonymous):

-2 | 1 3 4 8 | -2 -4 -------------------------- 1 1 2 4 is that right?

jimthompson5910 (jim_thompson5910):

close, the last row is correct

jimthompson5910 (jim_thompson5910):

the second row is not correct

OpenStudy (anonymous):

but you put that down yourself....

jimthompson5910 (jim_thompson5910):

I put this down -2 | 1 3 4 8 | -2 -2 ---------------------------------------- 1 1 but I wasn't finished (since I wanted to see if you could finish or not)

jimthompson5910 (jim_thompson5910):

The full and completed synthetic division table should be this -2 | 1 3 4 8 | -2 -2 -4 ---------------------------------------- 1 1 2 4

OpenStudy (anonymous):

that is basically what I put down. I just filled in the blanks...

jimthompson5910 (jim_thompson5910):

you were missing a -2 in the second row though

OpenStudy (anonymous):

okay. I get the point. Making this table still doesn't really give me the answer I'm looking for

jimthompson5910 (jim_thompson5910):

it does, if you know how to read it the bottom row is what is what we take away from the table the bottom row is 1, 1, 2, 4 The first three values form the coefficients of the quotient and the last value is the remainder

OpenStudy (anonymous):

got it

jimthompson5910 (jim_thompson5910):

so the coefficients of the quotient is 1, 1, 2 So the quotient is \[\Large x^2+x+2\]

OpenStudy (anonymous):

x^2+x+2 remainder is 4

jimthompson5910 (jim_thompson5910):

you nailed it

OpenStudy (anonymous):

thanks

jimthompson5910 (jim_thompson5910):

yw